OpenStudy (bekkah323):

You draw 2 cards from a well-shuffled deck of 52 cards. After you draw the first card, you do not replace it in the deck. What is the probability of P(get ace of spades on first draw and get queen of hearts on second draw)?

Directrix (directrix):

How many ace of spades cards are in the deck? Also, how many queen of hearts cards are in the deck? @bekkah323

Directrix (directrix):

I'm waiting for you to answer the questions above. @bekkah323

OpenStudy (bekkah323):

i think the answer is 1/51

Directrix (directrix):

This is the question I'd like to ask: How many ace of spades cards are in the deck? @bekkah323

OpenStudy (bekkah323):

1

Directrix (directrix):

Cool and also 1queen of hearts. Agree?

OpenStudy (bekkah323):

yes

Directrix (directrix):

The probability of getting the Ace of Spades is 1/52. The first card drawn is not replaced. There are now 51 cards in the deck. The probability of drawing the queen of hearts is 1/51. P( ace of spades on first draw and queen of hearts on second draw) = (1/52) * (1/51) = 1/2652 ---->